You are given the following ANOVA table, with n = 55. Compute the F statistic for the interaction term.

Source of
Variation

Sum of
Squares

Degrees of
Freedom

Mean
Square

F

Factor A

127

c-1 = 3

Factor B

784

r-1 = 7

Interaction

253

Error

5761

Total

12.0476

3.608

1.611

0.048

none of the above
Answer #1

Minitab Untitled - [Session] 日 | 盾范 勺 | t ↓鈍鷓 O@E|||幻园倫0E] I File Edit Data Calc Stat Graph Editor Tools Window Help Assistant 04-11-2018 01:04:44 Welcome to Minitab, press F1 for help. MTB > let k1 = 55-1 MTB > print kil Data Display K1 MTB > let k2 =54 _ (3 + 7 + 21) 54.0000 . .. df (Total) MTB > print k2 Data Display K2 MTB > let k3 = 127 + 784 + 253 + 5761 23.0000 .. . df (Error) MTB > print k3 Data Display K3 6925.00 .. . SS (Total) MTB > Let Mean Square = Sum of Squares/ Degrees of Freedom MTB > let k412.048 250.478 MTB > print k4 Data Display K4 0.0481000 .. . F statistic MTB >

F \ statistic = \frac{MS(Interaction)}{MS(Error)} = \frac{12.048}{250.478} = 0.0481

ans-> d) 0.048